OpenStudy (anonymous):

Convert the metric conversions. 13g = how many cg 0.6 m = how many mm 5.9 L = how many cL 49.74 W=how many mW 4000 cG= how many g 1576 mL = how many L 352 cm = how many m 26.8 mW = how many W

OpenStudy (anonymous):

It doesn't have cg

OpenStudy (anonymous):

medal for answers!

ganeshie8 (ganeshie8):

\(13 g = 13 g \times \frac{100\ cg}{1\ g} \)

OpenStudy (anonymous):

1300cg?

ganeshie8 (ganeshie8):

correct !

ganeshie8 (ganeshie8):

for other problems also, we just multiply wid a ratio that equals to 1 in a useful way

ganeshie8 (ganeshie8):

\(0.6 m = 0.6m\times \frac{1000\ mm}{1\ m}\)

OpenStudy (anonymous):

0.0006?

ganeshie8 (ganeshie8):

knw hw to multiply ha

ganeshie8 (ganeshie8):

\(0.6 m = 0.6m\times \frac{1000\ mm}{1\ m} = 0.6 \times 1000 mm = ?\)

OpenStudy (anonymous):

Oh LOL 600!

ganeshie8 (ganeshie8):

:)

ganeshie8 (ganeshie8):

see if u can try others on ur own

OpenStudy (anonymous):

5.9= 5900 49.74=49740 4000=40 1576=157600 352=35200 26.8=2680

ganeshie8 (ganeshie8):

first one is wrong, go step by step and slowly

ganeshie8 (ganeshie8):

\(5.9 L = 5.9 \times \frac{100\ cL}{1 \ L} = ?\)

OpenStudy (anonymous):

590?

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

Yay! What about the others?

ganeshie8 (ganeshie8):

its confusing to take everything at once, lets go one by one, tell me which one u r not sure of ?

OpenStudy (anonymous):

How about 49.74= 4970?

ganeshie8 (ganeshie8):

49.74 what= 4970 what

ganeshie8 (ganeshie8):

u must put units also, simply saying 49.74= 4970 is a crime !

OpenStudy (anonymous):

49.74 W = 4970 mW so sorry...

ganeshie8 (ganeshie8):

lol ok 4 missing 49.74 W = 49740 mW \(\color{green}{\checkmark}\)

OpenStudy (anonymous):

Yay! Here is the next one. 4000cg = 40 g

ganeshie8 (ganeshie8):

\(\color{green}{\checkmark}\)

OpenStudy (anonymous):

Yay! Next one. 1576 mL= 157600 L

OpenStudy (anonymous):

Yay! 352cm= 35200m